# Flyby-induced displacement effect: an analytic solution

### v3

@@ Abstract
We describe analytically the scattering of particles by a sandwich gravitational wave generated during a flyby. The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method. Our results confirm the prediction of Zel’dovich and Polnarev about certain “magical” amplitudes of the potential.
@@ I Introduction
Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories. Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : …although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.
@@
${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,.$ (2)
as profile.
The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
@@
In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal. It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt1 .,
@@
Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above. Our clue is that Eqn.(5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its (non-normalizable) zero-energy ground state solutions DM-1 .
@@ II Analytic solution for flyby
We first recall how things go for the simple Pöschl-Teller case. DM-1 ; DM-2 . Putting $t=\tanh U$, into eqn. (5) with profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ yields
$(1-t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}-2t\frac{dX^{\pm}}{dt}\mp\frac{g}{4}X^{\pm}=0\,,$ (7)
Let us consider first one of the components, say $X^{+}$. Then DM is obtained in terms of Legendre polynomials DM-1 when the coefficient in the linear-in-$X^{+}$ term is
@@
However passing to the other component, the linear-in-$X^{-}$ term has opposite sign, therefore DM would be obtained for $-g$. But this is impossible because $g$ has already been fixed. In conclusion, one has what we called “half DM” DM-2 : we get a DM solution in one or in the other sector, but not for both in general — unless the “wrong” component is set to zero.
Solutions for the Scarf potential ${\cal A}^{Scarf}$ (and similarly for derived Pöschl-Teller , ${\cal A}^{dPT}$), DM-2 ), are obtained by putting $t=\sinh{U}$. Then (5) is written
@@
The two equations here are again uncoupled and can therefore be solved independently. Note however the coefficient $t$ in front of the linear-in-$X^{\pm}$ term which compensates the $\pm$ sign change between the components. Therefore the same $g$ works for both components, as it will be seen in sec. III.
Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller . For derived Pöschl-Teller it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions in eqn. # (IV.16) of DM-2 .
Turning to Scarf, we note that (9) is of the generalised hypergeometric type which can be solved analytically by the Nikiforov-Uvarov (NU) method NikiUvar . Since a fairly detailed description of this simple yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , here we just list the main result. The algorithm, whose details are discussed in sec.III, yields the desired “magic” coefficient $g$ in (9),
@@
where $n$ is an integer. Then the trajectory is,
@@
$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,}\;\;$ (12)
@@
The analytic solutions (11) reproduce almost identically those obtained numerically in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case. For $n\geq 5$, though, the numerical procedure becomes tedious as it requires to calculate a high number of decimals. For high wave numbers the result becomes also inaccurate because the procedure used in DM-2 is only an approximation. Plotting instead the analytic solution (11)-(12), FIG.3, is straightforward, underlining the advantage of the analytical solution.
@@
Figure 3: DM geodesics for the Scarf profile (4) with wave numbers ${\bf n=5}$ and ${\bf n=6}$.
@@ III Transverse geodesics by the Nikiforov-Uvarov method
To have real solutions, we assume $k>0$, that is we choose th upper sign $k=\frac{\sqrt{1+16g^{2}}-1}{8}$ in Eq.(16). On the other side, the quantization condition Eq. (36) requires
@@
The $y_{n}(t)$ can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$. The first three of them are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$
It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. We have,
@@
can be easily integrated to yield (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
@@
$X^{\pm}(-t)=(-1)^{n}X^{\mp}(t)\,$ (23)
mentioned in sec.II. This curious pairing of the components is reminiscent of and is indeed closely related to supersymmetry MSUSY .
@@ IV Conclusion
The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.3, extending the observation made in DM-2 for the Gaussian and Pöschl-Teller profiles. The origin of half-waves is clear: their number is that of the zeros of the polynomial $y_{n}(t)$.
@@
Physical applications of various incarnations of the Scarf potential are reviewed in ref. Raposo . They include supersymmetry Scarfpot ; Scrf1 ; Compean ; Gendenshtein ; MSUSY and non-central potentials Dutt . The recent paper MSUSY , which is directly relevant for our purposes here, highlights the prominent role of SUSY for gravitational waves with Scarf profile. See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for further related work.

### v4

@@ Abstract
We describe analytically the scattering of particles by a sandwich gravitational wave generated during a flyby. The derivative-of-the-Gaussian profile proposed by Gibbons and Hawking is approximated by the hyperbolic Scarf potential, which allows for an exact analytic solution via the Nikiforov-Uvarov method. Our results confirm the prediction of Zel’dovich and Polnarev when the wave amplitude takes certain “magical” values.
@@ I Introduction
Initial studies Ehlers considered the Velocity Memory (VM) effect : after the wave has passed, the particles would fly apart with constant velocity along straight trajectories. Zel’dovich and Polnarev ZelPol advocated instead in favor of flyby: gravitational waves generated by a cluster of superdense stars (neutron stars or black holes) would yield the Displacement Memory (DM) effect : “…although the distance between a pair of two noninteracting bodies (such as satellites) will change, their relative velocity will become vanishingly small as the flyby concludes.”
@@
${\cal A}^{dG}\propto\frac{\;d}{dU}\Big{(}\exp[-U^{2}]\Big{)}\,$ (2)
as profile. The geodesic equations for (2) do not support any analytic solution, the profile can however be approximated by the derived Pöschl-Teller (dPT) potential PTeller ; Chakra ; DM-1 ,
@@
In this note we present a third approximation of the flyby profile (2), namely by the hyperbolic Scarf potential Scarfpot  [^4] The trigonometric Scarf potential was first introduced in Scarf to mimic the potential of a real crystal. It turned out to be the tip of the iceberg of a whole class of potentials with interesting properties, ranging from classical and quantum integrability Perelomov ; Perelomov1 to supersymmetric quantum mechanics Gendenshtein ; Natanson ; Dutt86 ; Levai ; Scrf1 . ,
@@
Our main result then says that for the profile (4) analytic solutions can be found which confirm the prediction of Zel’dovich and Polnarev ZelPol , supplemented by the quantization condition mentioned above. Our clue is that the geodesic equation (5) can be viewed as a time-independent Schrödinger equation and the DM trajectories correspond to its real (non-normalizable) zero-energy ground states DM-1 .
@@ II Analytic solution for flyby
We first recall how things go for the simple Pöschl-Teller case DM-1 ; DM-2 . Putting $t=\tanh U$ into eqn. (5) with profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ yields,
$(1-t^{2})\frac{d^{2}X^{\pm}}{dt^{2}}-2t\frac{dX^{\pm}}{dt}\mp\frac{g}{4}X^{\pm}=0\,.$ (7)
Let us consider first one of the components, say $X^{+}$. Then DM is obtained in terms of Legendre polynomials DM-1 when the coefficient in the linear-in-$X^{+}$ term is quantized, as
@@
However passing to the other component, the last, linear-in-$X^{-}$ term has opposite sign, therefore DM would be obtained for $-g$. But this is impossible because $g$ has already been fixed. In conclusion, one has what we called “half DM” DM-2 : we get a DM solution in one or in the other sector, but not for both in general — unless the “wrong” component is set to zero.
Solutions for the Scarf potential ${\cal A}^{Scarf}$ (and similarly for derived Pöschl-Teller , ${\cal A}^{dPT}$, DM-2 ), are obtained by putting $t=\sinh{U}$. Then (5) is written
@@
The two equations here are uncoupled and can therefore be solved independently. Note however the coefficient $t$ in front of the linear-in-$X^{\pm}$ term whose sign change compensates the $\pm$ sign change between the components. Therefore the same $g$ works for both components, as it will be seen in sec. III.
Finding the geodesics for dPT is considerably more difficult, though, as for “underived” Pöschl-Teller  ; it can be obtained in terms of confluent Heun functions DM-1 ; DM-2 with numerically calculated approximate expressions, see eqn. # (IV.16) of DM-2 .
Turning to Scarf, we note that (9) is of the generalised hypergeometric type which can be solved analytically by the Nikiforov-Uvarov (NU) method NikiUvar . Since a fairly detailed description of this simple yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , here we just list the main results. The algorithm, whose details are discussed in sec.III, yields the desired “magic” coefficient $g$ in (9),
@@
where $n$ is a natural integer. Then the trajectory is,
@@
$y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,$ (12)
@@
The analytic solutions (11) reproduce almost identically those obtained numerically in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case. For $n\geq 5$, though, the numerical procedure becomes tedious as it requires to calculate a high number of decimals. For high wave numbers the result becomes inaccurate because the procedure used in DM-2 is only an approximation. Plotting instead the analytic solution (11)-(12) yields accurate figures, underlining the advantage of the analytical approach.
@@ III Transverse geodesics by the Nikiforov-Uvarov method
To have real solutions, we assume $k>0$, that is we choose the upper sign $k=\frac{\sqrt{1+16g^{2}}-1}{8}$ in Eq.(16). On the other side, the quantization condition Eq. (37) requires
@@
There remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. We have,
@@
can be easily integrated to yield (up to an irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
@@
$X^{\pm}_{n}(-t)=(-1)^{n}X^{\mp}_{n}(t)\,$ (23)
mentioned in sec.II. Numerical evidence suggests that $t=0$ is not a root of $X^{\pm}_{n}$ neither for dPT nor for Scarf and, in fact, for a profile which is an order-odd derivative of an even function. On the contrary, it is a root when the profile is an order-even derivative.
The curious pairing of the components in (23) is reminiscent of and is indeed closely related to supersymmetry MSUSY .
The $y_{n}(t)$ can be expressed in terms of Routh-Romanovski polynomials Scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$. The first three of them are $y_{1}(t)\sim-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$
@@ IV Conclusion
The displacement effect (DM) arises when the wave zone contains $n$ half-waves as observed made DM-2 for the Gaussian and Pöschl-Teller profiles. The number of half-waves is equal to that of the zeros of the polynomial $y_{n}(t)$.
@@
Physical applications of various incarnations of the Scarf potential are reviewed in ref. Raposo . They include supersymmetry Scarfpot ; Scrf1 ; Compean ; Gendenshtein ; MSUSY and non-central potentials Dutt86 . The recent paper MSUSY , which is directly relevant for our purposes here, highlights the prominent role of SUSY QM for gravitational waves with Scarf profile. See also Benin ; Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for further related work.
@@
It has long been known (see, e.g., Gendenshtein ; Dutt86 ; Scarfpot ; Cooper94 ; Dutt92 ) that the Pöschl-Teller and (hyperbolic) Scarf potentials are parts of the same supersymmetric partner potentials, following from the shape invariant superpotential
$W(U)=A{\,}\tanh(U)+B{\,}{\rm sech}(U)\,.$ (25)
A recent paper MSUSY confirms our results using supersymmetric quantum mechanics with this superpotential.